A 180 μF defibrillator capacitor is charged to 1500 V. When fired through a patient's chest, it loses 95% of its charge in 40 ms. What is the resistance of the person's chest?
Answer #1

voltage on a capacitor, discharging   v = v₀e^(–t/τ)
v₀ is the initial voltage on the capacitor
v is the voltage after time t
R is resistance in ohms,
C is capacitance in farads
t is time in seconds
RC = τ = time constant

v = v₀e^(–t/τ)
0.05 = e^(–40m/τ)
(–40m/τ) = ln 0.05 = –3
τ = 40 ms/3 = 13.3 ms = RC
R = 13.3 m / 180 µ = 73.9 ohms